On the Standford-Binet test, the mean IQ is 100. A class of 21 kindergarten pupils were tested with a resulting mean of 108 and a standard deviation of 7.09. What is the lower bound of the 90% confidence interval of the average IQ of the kindergarten pupils?



1
Expert's answer
2022-02-28T16:08:17-0500

When the population standard deviation is unknown, the mean has a Student's t-distribution.

The critical value for "\alpha = 0.10" and "df = n-1 = 20" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.724718."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{X}+t_c\times\dfrac{s}{\sqrt{n}})""=(108-1.724718\times\dfrac{7.09}{\sqrt{21}},"

"108+1.724718\times\dfrac{7.09}{\sqrt{21}})"

"=(105.3318,110.6684)"

The lower bound of the 90% confidence interval of the average IQ of the kindergarten pupils is 105.3318.